300 liters of a 76% antifreeze solution had to be reduced to a 20% solution. How many liters of antifreeze should be extracted?
Answer by jojo14344(1513)

Ratio and Proportion amigo.
let "x"= amount in liters
so, (300/0.76) = (x/0.20)
x=(300*0.20) / 0.76
x=79 liters
therefore 300-79+=+221+liters to be Extracted
Thank you